Find the general solution of the given differential equation. y + 2y + y = 14e-t
Answer #1

SOLUTIONS (*) o Y+24+y=14 ēt complementary solution y+2y4y=0 » (D?+2D+I)y=0 [D= d Auxiliary Equation M2+2m+ICO (m+12=0 miaNow from yř +2 up + Yp. 14ět => Aēt_4042t+ A+ + 4Atét_204? Et+ Aézet 14ět QAēt = 14et JA=7 V V Hence Yp=7 tet Hence Complete